OpenStudy (vera_ewing):

Find the surface area of a cylinder with diameter 22 in and height 15 in.

OpenStudy (anonymous):

Surface Area of a Cylinder = 2 pi r^2 + 2 pi r h =2pi((22/2)^2)+2pi*(22/2)*15 =572pi =1796.99 =1797 inches Volume = pi*(r^2)*h =pi*(11^2)*15 =1815pi =5701.99 =5702inches^3

OpenStudy (muzzack):

Remember...to find the surface area of a cylinder use this formula: SA = 2•πi•r^2 +π•dh

OpenStudy (vera_ewing):

@dhiraj123 ty :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

can anyone help me with algebra 2???

OpenStudy (anonymous):

yes

OpenStudy (vera_ewing):

@sotoslugger sure

OpenStudy (anonymous):

just go to my page the question is there